# Towards Adversarial Robustness via Debiased High-Confidence Logit Alignment

### v1

@@ Conclusion
Evaluation Setup
Our evaluation consists of two main components: robustness performance and robust generalization assessment. To evaluate robustness performance, we use PGD (Madry et al. 2018), C&W (Carlini and Wagner 2017), and AutoAttack (AA) (Croce and Hein 2020b) within the $\ell_{\infty}$-norm. AutoAttack includes several attack methods such as APGD-DLR (Croce and Hein 2020b), APGD-CE (Croce and Hein 2020b), FAB (Croce and Hein 2020a), and Square (Andriushchenko et al. 2020). Adversarial attacks are generated using a step size of $\alpha=2/255$ and a specified maximum $\ell_{\infty}$-norm. Note that the “Clean” indicates natural examples unaffected by adversarial perturbations.
To evaluate robust generalization, we introduce the concept of the robust generalization gap (Yin et al. 2023) denoted as the “Robust Gap” quantifying the difference in robust performance between training and test sets under adversarial attacks. A smaller robust gap indicates improved robust generalization, reflecting reduced vulnerability to robust overfitting in the adversarial training model.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) MART (b) UIAT (c) SGLR (d) DHAT
Figure 9: Two-dimensional t-SNE embeddings of depth features extracted from the penultimate layer of the ResNet-18 model, trained on CIFAR-10 using various methods. The visualization illustrates the distinctiveness and compactness of the learned feature representations for each method.
Performance Under Transfer Attack
In this part, we evaluate our proposed model’s performance under transfer attacks and compare it with the UIAT (Dong et al. 2023). Transfer attacks evaluate a model’s robustness by testing its performance against adversarial examples generated from different source models, which simulate real-world conditions where attackers may use varied strategies. This testing is essential to ensure that a model’s adversarial defenses generalize effectively beyond its training environment and are robust against diverse attack methods. By evaluating the performance under transfer attacks, we ensure that our model is robust not only to attacks generated by its own architecture but also to those from different models, providing a more comprehensive measure of robustness.
Robustness Performance
Robustness Across Various Attack Types
Generalization Across Various Source Models
Robustness under Various $\epsilon$ Attacks
In this part, we evaluate the robustness of our proposed model under various levels of adversarial perturbations, quantified by different $\epsilon$ values. We compare our model with two baseline methods, MART (Wang et al. 2020) and UIAT (Dong et al. 2023), and the state-of-the-art SGLR method (Li et al. 2024b) across various attack types, including FGSM, PGD-10, PGD-20, and C&W. We aim to demonstrate the superior generalization and robustness of our model under progressively challenging adversarial conditions. Figure 8 shows the accuracy of different methods under various $\epsilon$ values for each attack type.
@@
Robustness Across Various Attack Types
@@
Generalization Against Various Attacks
@@
Visualization of Learned Feature Distributions
In this part, we provide a comparative analysis of the feature distributions learned by our proposed method, DHAT, and benchmark it against MART (Wang et al. 2020), UIAT (Dong et al. 2023), and SGLR (Li et al. 2024b). Figure 9 illustrates these feature distributions using t-SNE visualization, with embeddings obtained from the penultimate layer of the ResNet-18 model trained on CIFAR-10.
Our method, DHAT, demonstrates a marked improvement in feature representation. Specifically, the features learned by DHAT are more compact and distinct compared to those produced by the other methods. This enhanced discriminability stems from DHAT’s effectiveness in mitigating spurious correlation bias. By aligning adversarial examples with debiased high-confidence logits, DHAT facilitates the learning of features that are both robust and discriminative, leading to improved performance, particularly in margin-based attacks such as C&W attacks.
When compared between MART and UIAT, MART demonstrates superior feature distinguishability. This advantage stems from MART’s alignment of logit distributions based on natural examples, which promotes the learning of more distinguishable features. In contrast, UIAT aligns the logits of adversarial examples with those of inverse adversarial examples, which are correlated with biased feature activation, resulting in less distinguishable feature representations. Additionally, SGLR’s performance in feature distinguishability is inferior to that of DHAT. This is because SGLR relies on soft-label distributions during training, which tends to rely on spurious correlations, thereby impairing class differentiation.
In summary, DHAT demonstrates superior performance in t-SNE feature distributions, showcasing enhanced feature distinguishability and more effective representations. This improvement is attributed to DHAT’s ability to mitigate spurious correlation bias, which facilitates more robust and accurate feature representations.
Refer to caption Refer to caption
(a) UIAT (b) DHAT
Figure 10: The learning curves for natural accuracy (%) and adversarial accuracy (%) under PGD-10 attack of ResNet-18 on CIFAR-10 dataset.
Robust Overfitting Analysis
In this part, we compare the learning curves of DHAT and UIAT to analyze robust overfitting and generalization. Experiments were conducted using ResNet-18 on the CIFAR-10 dataset, with a total of 200 training epochs for better visualization. Learning rate decay was applied at epochs 100 and 150, reducing the learning rate to 0.1.
Figure 10 shows the learning curves for DHAT and UIAT. Our analysis shows that neither DHAT nor UIAT exhibits a significant or irreversible decline in robustness performance after learning rate adjustments (Yu et al. 2022). This indicates that both methods maintain consistent robustness throughout the training process.
However, DHAT consistently shows a smaller robust generalization gap and standard generalization gap between the training and test sets than UIAT. This reduced performance gap underscores DHAT’s superior generalization ability across different distributions, such as adversarial attack conditions or transitions to unseen test sets. The reduced performance disparity implies that DHAT effectively combats robust overfitting, leading to more reliable and consistent adversarial robustness. In contrast, UIAT shows a more pronounced performance decrease, indicating the degraded generalization capability.
In conclusion, these findings confirm that DHAT not only enhances adversarial robustness but also improves robust generalization, representing a significant advancement in adversarial training techniques.

### v2

@@ 7 Algorithm Details
8 Performance Under Transfer Attack
In this part, we evaluate our proposed model’s performance under transfer attacks and compare it with the UIAT [9]. Transfer attacks evaluate a model’s robustness by testing its performance against adversarial examples generated from different source models, which simulate real-world conditions where attackers may use varied strategies. This testing is essential to ensure that a model’s adversarial defenses generalize effectively beyond its training environment and are robust against diverse attack methods. By evaluating the performance under transfer attacks, we ensure that our model is robust not only to attacks generated by its own architecture but also to those from different models, providing a more comprehensive measure of robustness.
8.1 Robustness Performance
8.2 Robustness Across Various Attack Types
8.3 Generalization Across Various Source Models
Table 7: Comparison of robustness (%) and robust generalization gap (%) for models trained on generated data. The bolded numbers indicate the best performance.
CIFAR-10 ResNet-18 WRN28-10
[…]
MART [45] 83.45 49.45 2.91 84.26 51.95 8.73
AWP [49] 83.78 50.79 2.03 84.10 53.29 6.11
FSR [21] 83.19 50.53 2.35 83.88 53.03 7.05
CFA [47] 84.97 50.85 2.98 85.81 53.35 8.94
UIAT [9] 85.10 52.09 3.04 86.73 54.59 9.41
SGLR [27] 86.35 51.27 2.75 87.72 53.77 11.25
DHAT (Ours) 87.62 54.42 0.83 88.94 56.92 2.48
CIFAR-100 ResNet-18 WRN28-10
[…]
MART [45] 54.73 27.70 2.88 55.87 30.20 8.86
AWP [49] 57.55 29.33 2.37 59.71 31.83 7.13
FSR [21] 58.10 28.94 2.47 59.03 30.44 7.40
CFA [47] 60.13 28.85 3.01 61.56 29.61 9.13
UIAT [9] 59.92 28.48 4.47 60.24 30.98 14.32
SGLR [27] 61.25 29.10 4.15 62.39 30.15 17.05
DHAT (Ours) 63.14 32.21 0.98 64.80 34.71 2.96
9 Performance with Generated Data
We employ Diffusion Denoising Probabilistic Models (DDPM) [17] to generate an additional 50K samples for both CIFAR-10 and CIFAR-100 datasets following the [46]. This synthetic data is then used to augment the training of both our method and all baseline models. The performance comparisons, as shown in Table 7, illustrate the impact of additional data on robustness and generalization.
9.1 Robustness Performance
The results in Table 7 demonstrate that our method outperforms baseline methods in terms of robustness when trained with the additional synthetic data. This improvement highlights the effectiveness of our approach in leveraging extra data to enhance model robustness, particularly under adversarial conditions.
9.2 Generalization Performance
We observe that, compared to the results in Table 1, most models trained with the additional data exhibit a reduced Robust Gap, indicating improved generalization. However, both UIAT and SGLR exhibit limited improvements in robust generalization. These methods rely on spurious correlations during training, which hinders their generalization performance even with the added data. Although UIAT and SGLR benefit from enhanced robustness due to increased data diversity, their robust generalization remains suboptimal, likely due to their dependence on non-essential features. This finding underscores the unique advantage of our method in achieving both robust accuracy and generalization with enriched datasets.
Table 8: Comparison of robustness (%) and robust generalization gap (%) for using various attention map generation techniques using WRN28-10 on the CIFAR-10.
[…]
- 82.94 58.66 54.11 52.17 7.92
Grad-CAM [33] 83.95 60.49 55.27 53.10 3.51
Integrated-Grad [41] 83.97 60.35 55.18 53.04 3.68
SOLO [44] 84.26 61.60 56.74 54.93 3.09
SAM [22] 85.65 62.44 58.10 56.38 2.46
10 Impact of Foreground-Background Recognition Techniques
In this part, we investigate the influence of different attention map generation techniques on the performance of Debiased High-Confidence Logit Regularization (DHAT), as seen in Table 8. Our exploration extends beyond the initially employed Grad-CAM method to encompass a variety of attention map generation approaches, thereby providing a comprehensive analysis of their effects on model robustness and generalization.
@@ 10 Impact of Foreground-Background Recognition Techniques
10.1 Exploration of Various Attention Map
The primary approach detailed in the main text utilizes Grad-CAM [33], a widely adopted method in weakly supervised object segmentation [28], to extract foreground and background feature maps efficiently. Grad-CAM offers a balance between computational efficiency and effectiveness, making it suitable for large-scale evaluations. However, to understand the broader applicability and potential improvements, we incorporated additional attention map generation techniques, including Integrated Gradients [41], SOLO [44], and SAM [22].
10.2 Evaluation of Advanced Attention Map Models
Incorporating more sophisticated attention map generation models, such as SAM, demonstrated enhanced robustness and reduced robust generalization gaps compared to simpler methods like Grad-CAM. These advanced models provide finer and more accurate segmentation of foreground and background features, which in turn leads to better alignment of high-confidence logits under adversarial conditions. However, this improvement in performance comes at the cost of increased computational overhead. More complex models require greater processing power and longer training times, which may be a limiting factor in resource-constrained environments.
10.3 Trade-Off Between Performance and Computational Efficiency
While the adoption of advanced attention map techniques can yield superior performance metrics, researchers must consider the trade-offs involved. Enhanced models may offer marginal gains in robustness and generalization, but the additional computational resources and time required may not always justify these benefits, especially in applications where real-time processing is essential. Therefore, the choice of attention map generation technique should be informed by the specific requirements and constraints of the deployment scenario.
11 Robustness under Various Magnitude Attacks
In this part, we evaluate the robustness of our proposed model under various levels of adversarial perturbations, quantified by different $\epsilon$ values. We compare our model with two baseline methods, MART [45] and UIAT [9], and the state-of-the-art SGLR method [27] across various attack types, including FGSM, PGD-10, PGD-20, and C&W. We aim to demonstrate the superior generalization and robustness of our model under progressively challenging adversarial conditions. Figure 8 shows the accuracy of different methods under various $\epsilon$ values for each attack type.
@@ 11 Robustness under Various Magnitude Attacks
11.1 Robustness Across Various Attack Types
@@
11.2 Generalization Against Various Attacks